# SCLIP: Rethinking Self-Attention for Dense Vision-Language Inference

### v2

@@ Abstract
Recent advances in contrastive language-image pretraining (CLIP) have demonstrated strong capabilities in zero-shot classification by aligning visual representations with target text embeddings in an image level. However, in dense prediction tasks, CLIP often struggles to localize visual features within an image and fails to give accurate pixel-level predictions, which prevents it from functioning as a generalized visual foundation model. In this work, we aim to enhance CLIP’s potential for semantic segmentation with minimal modifications to its pretrained models. By rethinking self-attention, we surprisingly find that CLIP can adapt to dense prediction tasks by simply introducing a novel Correlative Self-Attention (CSA) mechanism. Specifically, we replace the traditional self-attention block of CLIP vision encoder’s last layer by our CSA module and reuse its pretrained projection matrices of query, key, and value, leading to a training-free adaptation approach for CLIP’s zero-shot semantic segmentation. Extensive experiments show the advantage of CSA: we obtain a 38.2% average zero-shot mIoU across eight semantic segmentation benchmarks highlighted in this paper, significantly outperforming the existing SoTA’s 33.9% and the vanilla CLIP’s 14.1%. Code is available at https://github.com/wangf3014/SCLIP.
Figure 1: Open-vocabulary semantic segmentation examples. In each triplet, Upper left: original image; Lower left: vanilla CLIP’s segmentation result; and Right: our segmentation result. We evaluate on two images from COCO [5] (the macaw and vegetable examples) and four high-resolution in-the-wild images, where our SACLIP consistently generates high quality segmentation masks, yet vanilla CLIP fails to correctly localize the objects. We display the corresponding text query of each segmentation mask, where “g. retriever” and “b. collie” in the first example denote golden retriever and border collie, respectively.
@@ 1 Introduction
Specifically, CLIP performs zero-shot classification by matching image-level representations with a range of target text embeddings, by which it achieves over 70% test accuracy on ImageNet [13] when paired with proper prompting strategies [42]. However, directly transferring this inference protocol to semantic segmentation fails to achieve favorable results. For example, when equipped with a ViT-Base/16 [17] encoder and fed with a 224$\times$224 resolution input image, CLIP can obtain a 14$\times$14 dense feature map; and by simply associating such patch-level representations with text embeddings, CLIP yields a mere 3.1% mIoU on ADE20k [65] and 5.7% mIoU on COCO-Stuff [5]. Considering the supervised counterparts that often produce around 40% mIoU on this two benchmarks, this result is not really comparable. As a result, CLIP still relies on very careful finetuning and in-domain adaptations for downstream dense prediction tasks [37, 68, 60].
In this work, we investigate in CLIP’s potential for dense prediction and find out whether the weak supervisions of CLIP can benefit various vision tasks with minimal downstream adaptations. We start with a qualitative analysis. As shown in Figure 1, we conduct simple open-vocabulary semantic segmentation experiments on six sample images from COCO [5] or in the wild, where the vanilla CLIP model often presents incorrect dense predictions and noisy segmentation masks (lower left for each triplet). However, despite its poor semantic segmentation performance, we find that CLIP is actually able to roughly recognize what objects appear in the image yet mostly wrongly localizes them. For instance, in the fifth example of Figure 1, we set 10 target categories including flamingo, water, land, with distractors such as sky, building, and person, but although CLIP accurately obtains the correct categories such as water and flamingo, it predicts the opposite localizations (i.e., predicts water for flamingos and flamingo for water and land).
Figure 2: Final layer attention maps of vanilla CLIP with a ViT-Base/16 image encoder. We display four attention maps for each example, and the source locations are marked in different colors in the image. It shows that each local visual token attends to a wide range of positions and the attention maps often share similar patterns, which indicates CLIP learns spatial-invariant visual features, i.e., the local features are invariant to their spatial positions.
@@
Empirical study on our SCLIP model showcases its notable effectiveness, with yielding both impressive qualitative and quantitative outcomes: we obtain an average mIoU of 38.2% over eight semantic segmentation benchmarks such as PASCAL Context [38] and COCO-Stuff [5], substantially outperforming the existing state-of-the-art methods such as MaskCLIP [66] (30.3%), GroupViT [57] (30.7%), and TCL [8] (33.9%) that support zero-shot and open-vocabulary semantic segmentation. In Figure 1, we also show the qualitative results obtained by SCLIP for images in the COCO [5] dataset and in the wild, where our model yields very clear and accurate segmentation masks, especially for the high-resolution inputs (e.g., the case of two dogs sitting on the boat).
The primary contributions of this work can be summarized as follows. First, We identify the reasons of vanilla CLIP’s failure in semantic segmentation, and address them by introducing a novel Correlative Self-Attention (CSA) mechanism, while extensive experiments demonstrate significant results. Next, our SCLIP approach outperforms the existing methods [66, 57, 47, 8] with neither fine-tuning nor any additional parameters given a pretrained CLIP model, which validates the good transferability of vision-language models in dense prediction tasks. Further, in this work, a minimal modification to CLIP yields very significant improvements in semantic segmentation, which provides with an important data point that the weakly-supervised pretraining paradigm with language guidance has very good potentials to function as a visual foundation model that supports a wide range of downstream tasks.
@@ 2 Related Work
Transferable Visual Foundation Models. Self-supervised pretraining has recently demonstrated good potentials in learning transferable visual representations. The models pretrained with re-constructive objectives such as masked image modeling [25, 2, 55] or discriminative objectives such as contrastive learning [24, 9, 21, 7] exhibit strong capabilities in adapting various visual tasks when sufficient downstream training data is available. Similarly, denoising diffusion models [48, 15, 26] that allow high-resolution conditional image generation and Segment Anything models [29, 71] that facilitate semantic-agnostic image segmentation can also serve as foundation models with transferable visual features.
@@
Open-Vocabulary Segmentation. To fully utilize the advancements of vision-language models in zero-shot and open-vocabulary visual inference, extensive follow-up work has been initiated to investigate their applications in dense prediction tasks. For example, GroupViT [57] introduces group tokens into its vision encoder and pretrains with language guidance, leading to an open-vocabulary model that well applies to semantic segmentation tasks. Also, MaskCLIP [66] makes simple modification to vision transformers and enables CLIP’s coarse feature localization. The study of language-guided segmentation is continuously explored [20, 59, 37, 68, 60, 58, 8, 34, 47, 44, 31, 64, 61, 39].
Self-Attention for Dense Visual Features. A series of related research has demonstrated that the potential of vision transformers in extracting dense visual features can be augmented by employing varied self-attention mechanisms. For example, in contrast to the vanilla self attention used in CLIP and conventional vision transformers [42, 17], the Local Attention mechanism constrains the spatial feature aggregation within a local window so that to encourage fine-grained features [35, 36, 69, 54, 50]. MaskCLIP discards the processing of query and key vectors in its last transformer layer, which can also be considered as a specific case of local attention with window size set to one [66]. In addition, some segmentation or detection-oriented transformer models leverage cross attention to map local visual features into semantic tokens [57, 10, 11, 6, 49]. Also, the models equipped with Axial Attention [52, 53, 16] or Deformable Attention [70, 56] demonstrate strong capabilities in dense prediction.
@@ 3.3 Segmentation-Adapted CLIP Model
To develop our SCLIP approach, we employ a pretrained CLIP model with a ViT-Base/16 [17] image encoder as backbone. Generally, when it comes to adapting CLIP into a downstream task without introducing additional parameters, we actually regard its last or last several layers as a task-specific decoder head. Following MaskCLIP [66], we regard the last transformer block of CLIP’s image encoder as the decoding layer to implement the adaptations while leaving the remaining components unchanged.
@@ 4 Experiments
VOC21 Context60 COCO-Obj. VOC20 City. Context59 ADE COCO-Stf.
@@
$*{}^{*}$ 19.8 8.7 10.4 54.2 7.0 11.7 3.6 5.9 15.2
$*{}^{*}$ 52.0 28.2 22.6 72.1 30.1 31.5 14.0 20.0 33.8
$*{}^{*}$ 52.7 19.5 27.9 81.5 21.7 24.4 11.8 16.9 32.1
$*{}^{*}$ 27.2 21.9 17.3 62.4 23.2 24.7 12.4 16.3 25.7
$*{}^{*}$ 55.0 30.4 31.6 83.2 24.3 33.9 17.1 22.4 37.2
$*{}^{*}$ (ours) 61.7 31.5 32.1 83.5 34.1 36.1 17.8 23.9 40.1
@@ 4.1 Experiment Settings
Datasets. We evaluate our method on six commonly used semantic segmentation benchmarks, including PASCAL VOC 2012 [18], PASCAL Context [38], Cityscapes [12], ADE20k [65], COCO-Stuff and COCO-Object [5]. Considering the background category, we additionally evaluate on two variant datasets for PASCAL VOC and PASCAL Context. For clear reference, we denote VOC21, Context60 as the original datasets with a background class, and VOC20, Context59 as the variant without this category. In the existing works such as GroupViT [57] and TCL [8], they evaluate with input images resized to have a shorter side of 448 and then performing slide inference with a 448$\times$448 window and 224 stride. However, in our experiments, we find that using a smaller input size with denser sliding stride can lead to slightly higher results (e.g., 0.2% mIoU on PASCAL Context). Specifically, we resize input images with a short side of 336 and perform slide inference with a 224$\times$224 window and 112 stride. This protocol introduces a similar level of computation as that of GroupViT, yet better fits the original input size of CLIP (e.g., 224 for ViT-Base) and is also friendly to parallel computing. For Cityscapes [12], we resize with a 560 shorter side due to the particular high resolution of its original images. A detailed comparison of image pre-processing protocols can be found in Table 4.
@@ 4.3 Ablation Study
Mode VOC21 Context COCO-Stf.
@@
Mode Img size Window Stride Flops mIoU
#1 224 224 112 1×1$\times$1 × 56.5
#2 336 224 112 4$\sim$ 4$\times$ 4 × 59.1
#3 336 336 112 3$\sim$ 3$\times$ 3 × 58.6
#4 448 224 112 9$\sim$ 9$\times$ 9 × 60.4
#5 448 448 224 4$\sim$ 4$\times$ 4 × 58.9

### v4

@@ Abstract
Recent advances in contrastive language-image pretraining (CLIP) have demonstrated strong capabilities in zero-shot classification by aligning visual and textual features at an image level. However, in dense prediction tasks, CLIP often struggles to localize visual features within an image and fails to attain favorable pixel-level segmentation results. In this work, we investigate in CLIP’s spatial reasoning mechanism and identify that its failure of dense prediction is caused by a location misalignment issue in the self-attention process. Based on this observation, we propose a training-free adaptation approach for CLIP’s semantic segmentation, which only introduces a very simple modification to CLIP but can effectively address the issue of location misalignment. Specifically, we reform the self-attention mechanism with leveraging query-to-query and key-to-key similarity to determine attention scores. Remarkably, this minimal modification to CLIP significantly enhances its capability in dense prediction, improving the original CLIP’s 14.1% average zero-shot mIoU over eight semantic segmentation benchmarks to 38.2%, and outperforming the existing SoTA’s 33.9% by a large margin. Code is available at https://github.com/wangf3014/SCLIP.
Keywords:
CLIP Self-attention Semantic segmentation
@@ 1 Introduction
Figure 1: Open-vocabulary semantic segmentation examples. We evaluate on two images from COCO [5] (the 3rd and the 5th examples) and three high-resolution images in the wild, where our SCLIP consistently generates high quality segmentation masks yet the original CLIP fails to correctly localize objects. We display the corresponding text query of each segmentation mask, where “g. retriever” and “b. collie” in the first example denote golden retriever and border collie, respectively.
@@
Specifically, CLIP performs zero-shot classification by matching image-level representations with a range of target text embeddings, by which it achieves over 70% test accuracy on ImageNet [13] when paired with proper prompting strategies [44]. However, directly transferring this inference protocol to semantic segmentation fails to achieve favorable results. For example, when equipped with a ViT-Base/16 [17] encoder and fed with a 224$\times$224 resolution input image, CLIP can obtain a 14$\times$14 dense feature map; and by simply associating such patch-level representations with text embeddings, CLIP yields a mere 3.1% mIoU on ADE20k [70] and 5.7% mIoU on COCO-Stuff [5]. Considering the supervised counterparts that often produce around 40% mIoU on this two benchmarks, this result is not really comparable. As a result, CLIP still relies on careful finetuning and in-domain adaptations for downstream dense prediction tasks [39, 73, 64].
In this work, we investigate in CLIP’s potential for dense prediction and find out whether the weak supervisions of CLIP can benefit various vision tasks with minimal downstream adaptations. We start with a qualitative analysis. As shown in Figure 1, we conduct simple open-vocabulary semantic segmentation experiments on five sample images from COCO [5] or in the wild, where the vanilla CLIP model often presents incorrect dense predictions and noisy segmentation masks. However, despite its poor semantic segmentation performance, we find that CLIP is actually able to roughly recognize what objects appear in the image yet wrongly localizes them. For instance, in the second example, we set 10 target categories including flamingo, water, land, with distractors such as sky, building, and person, but although CLIP accurately obtains the correct categories such as water and flamingo, it predicts the opposite localizations (i.e., predicts water for flamingos and flamingo for water and land).
Figure 2: Final layer attention maps of vanilla CLIP with a ViT-Base/16 image encoder. We display the attention maps of four points (marked in different colors) for each example. It shows that each local visual token attends to a wide range of positions and the attention maps often share similar patterns, indicating that CLIP learns spatial-invariant visual features.
@@
Empirical study on our SCLIP model showcases its notable effectiveness, with yielding both impressive qualitative and quantitative outcomes: we obtain an average mIoU of 38.2% over eight semantic segmentation benchmarks such as PASCAL Context [40] and COCO-Stuff [5], substantially outperforming the existing state-of-the-art methods such as MaskCLIP [71] (30.3%), GroupViT [61] (30.7%), and TCL [8] (33.9%) that support zero-shot and open-vocabulary semantic segmentation. In Figure 1, we also show the qualitative results obtained by SCLIP for images in the COCO [5] dataset and in the wild, where our model yields very clear and accurate segmentation masks, especially for the high-resolution inputs (e.g., the case of two dogs sitting on the boat). The primary contributions of this work can be summarized as follows:
  • First, We identify the reasons of CLIP’s failure in semantic segmentation, and address them by introducing a novel Correlative Self-Attention (CSA) mechanism, while extensive experiments demonstrate significant results.
  • Next, our SCLIP approach outperforms the existing methods [71, 61, 49, 8] with neither fine-tuning nor any additional parameters given a pretrained CLIP model, which validates the good transferability of vision-language models in dense prediction tasks.
  • Further, in this work, a minimal modification to CLIP yields very significant improvements in semantic segmentation, which provides with an important data point that the weakly-supervised pretraining paradigm with language guidance has very good potentials to function as a visual foundation model that supports a wide range of downstream tasks.
@@ 2 Related Work
Transferable Visual Foundation Models. Self-supervised pretraining has recently demonstrated good potentials in learning transferable visual representations. The models pretrained with re-constructive objectives such as masked image modeling [23, 2, 58] or discriminative objectives such as contrastive learning [24, 9, 21, 7, 26, 54] exhibit strong capabilities in adapting various visual tasks when sufficient downstream training data is available. Similarly, denoising diffusion models [50, 15, 27] that allow high-resolution conditional image generation and Segment Anything models [30, 76] that facilitate semantic-agnostic image segmentation can also serve as foundation models with transferable visual features.
@@
Open-Vocabulary Segmentation. To fully utilize the advancements of vision-language models in zero-shot and open-vocabulary visual inference, extensive follow-up work has been initiated to investigate their applications in dense prediction tasks. For example, GroupViT [61] introduces group tokens into its vision encoder and pretrains with language guidance, leading to an open-vocabulary model that well applies to semantic segmentation tasks. Also, MaskCLIP [71] and CLIP Surgery [33] make simple modification to vision transformers and enables CLIP’s coarse feature localization. The study of language-guided segmentation is continuously explored [20, 62, 39, 73, 64, 63, 8, 36, 49, 46, 32, 69, 65, 41, 59].
Self-Attention for Dense Visual Features. A series of related research has demonstrated that the potential of vision transformers in extracting dense visual features can be augmented by employing varied self-attention mechanisms. For example, in contrast to the vanilla self attention used in CLIP and conventional vision transformers [44, 17], the Local Attention mechanism constrains the spatial feature aggregation within a local window so that to encourage fine-grained features [38, 37, 74, 57, 52]. Localized visual features can also be encouraged with modified self-attention mechanisms such as MaskCLIP [71] which discards the processing of query and key vectors in its last transformer layer (equivalent to local attention with window being one) and MSSA [67] that reduces the attention projections into a single matrix. In addition, some segmentation or detection-oriented transformer models leverage cross attention to map local visual features into semantic tokens [61, 11, 10, 6, 51]. Also, the models equipped with Axial Attention [56, 55, 16] or Deformable Attention [75, 60] demonstrate strong capabilities in dense prediction.
@@ 3.3 Segmentation-Adapted CLIP Model
To develop our SCLIP approach, we employ a pretrained CLIP model with a ViT-Base/16 [17] image encoder as backbone. Generally, when it comes to adapting CLIP into a downstream task without introducing additional parameters, we actually regard its last or last several layers as a task-specific decoder head. Based on our observation of the self-attention patterns in different layers, we regard the last transformer block of CLIP’s image encoder as the decoding layer to implement the adaptations while leaving the remaining components unchanged.
@@ 4 Experiments
VOC21 Context60 Object VOC20 City. Ctx59 ADE20k Stuff.
@@
CLIP 19.8 8.7 10.4 54.2 7.0 11.7 3.6 5.9 15.2
MaskCLIP 52.0 28.2 22.6 72.1 30.1 31.5 14.0 20.0 33.8
GroupViT 52.7 19.5 27.9 81.5 21.7 24.4 11.8 16.9 32.1
ReCo 27.2 21.9 17.3 62.4 23.2 24.7 12.4 16.3 25.7
TCL 55.0 30.4 31.6 83.2 24.3 33.9 17.1 22.4 37.2
SCLIP (ours) 61.7 31.5 32.1 83.5 34.1 36.1 17.8 23.9 40.1
@@ 4.1 Experiment Settings
Datasets. We evaluate our method on six commonly used semantic segmentation benchmarks, including PASCAL VOC 2012 [18], PASCAL Context [40], Cityscapes [12], ADE20k [70], COCO-Stuff and COCO-Object [5]. Considering the background category, we additionally evaluate on two variant datasets for PASCAL VOC and PASCAL Context. For clear reference, we denote VOC21, Context60 as the original datasets with a background class, and VOC20, Context59 as the variant without this category. In prior works such as GroupViT [61] and TCL [8], they evaluate with input images resized to have a shorter side of 448 and then performing slide inference with a 448$\times$448 window and 224 stride. However, in our experiments, we find that using a smaller input size with denser sliding stride can lead to slightly higher results (e.g., 0.2% mIoU on PASCAL Context). Specifically, we resize input images with a short side of 336 and perform slide inference with a 224$\times$224 window and 112 stride. This protocol introduces a similar level of computation as that of GroupViT, yet better fits the original input size of CLIP (e.g., 224 for ViT-Base) and is also friendly to parallel computing. For Cityscapes [12], we resize with a 560 shorter side due to the particular high resolution of its original images. A detailed comparison of image pre-processing protocols can be found in Table 4.
@@ 4.3 Ablation Study
Mode PASCAL VOC PASCAL Context COCO-Stuff
@@
Mode  Img size  Window  Stride  Flops  VOC21 Context59 COCO-Stuff
#1 224 224 112 1×1$\times$1 × 56.5 32.0 20.5
#2 336 224 112 4$\sim$ 4$\times$ 4 × 59.1 34.2 22.4
#3 336 336 112 3$\sim$ 3$\times$ 3 × 58.6 34.1 21.9
#4 448 224 112 9$\sim$ 9$\times$ 9 × 60.4 35.3 23.4
#5 448 448 224 4$\sim$ 4$\times$ 4 × 58.9 34.3 22.1
@@ Acknowledgements
Figure 5: Additional visualization results on PASCAL VOC. “GT” denotes ground truth.
Figure 6: Additional visualization results on COCO-Object. “GT” denotes ground truth.